OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x-8y=0 -3x+8y=-7 A.INFINITELY MANY B.2 C.1<--I THINK D.0 ANYONE PLEASE HELP WILL GIVE MEDAL THANKS

OpenStudy (anonymous):

@E>Nattie12345

OpenStudy (anonymous):

@IAMBATMAN

OpenStudy (anonymous):

@E>Nattie12345

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@mar

OpenStudy (anonymous):

anyone please

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@PookieBear555

Directrix (directrix):

If you use the Addition method: 4x-8y=0 -3x+8y=-7 ----------- x + 0 = -7 x = -7 You can find the y by substitution of -7 for x in either equation.

OpenStudy (anonymous):

im sorry so is it 0 solutions

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

So, which of these does that indicate: A.INFINITELY MANY B.2 C.1<--I THINK D.0

Directrix (directrix):

If cannot be 0 solutions because we got a solution with x value -7

OpenStudy (anonymous):

so im right :,)

Directrix (directrix):

Correct. One solution.

OpenStudy (anonymous):

i goingt ot give you a medal for being nice

Directrix (directrix):

Thank you.

OpenStudy (anonymous):

no problem can you help me on some more please

Directrix (directrix):

I will try. Close this thread and post the next question in a new thread.

OpenStudy (anonymous):

hold on im writing the problem down so i could type it in

OpenStudy (anonymous):

ok